# Diversity Methods for Improving Convergence and Accuracy of Quantum Error Correction Decoders Through Hardware Emulation

### v1

@@ 2.0.5 Control layer
This communication is particularly crucial when the number of noise generators $NG$ is less than the number of required noise samples $n$. In such cases, $\left\lceil\frac{NG}{n}\right\rceil$ cycles are necessary, and the decoder may finish before the next round of samples is generated due to the early stopping criterion (see Scenario B in Fig. 2).
@@
Counterintuitively, in regions where the physical error rate is reduced, the time bottleneck can occur during noise generation. This is because the decoders require a large amount of area resources and leave insufficient space on the FPGA to introduce enough noise generators to perform all the computations in parallel.
@@
We can generally categorize two scenarios: one in which the decoder consumes more clock cycles than the noise generation (Scenario A in Fig. 2), and another where the decoder is faster (Scenario B). For instance, in a fully parallel BP decoder, the dividing line between both scenarios occurs when the number of iterations required for convergence is fewer than $\left\lceil\frac{NG}{2n}\right\rceil$, assuming two clock cycles per iteration, one for computing the check nodes and another for computing the variable nodes.
@@
The signals exchanged between both finite state machines include “noise ready," “generate noise," and “end of decoding." For simplicity, other control signals, such as those used to store error patterns in RAM when a logical failure occurs (for subsequent a posteriori analysis), have been omitted.
3 Results for BP decoders
The emulator discussed in the previous section was tested with various QLDPC codes, which are defined as ($n$, $k$, $d$), where $n$ is the number of physical qubits, $k$ is the number of logical qubits, and $d$ is the minimum distance of the code. Fig. 3 presents a small sample of eight codes with check node degrees of six and eight, decoded using the scaled min-sum (MS) algorithm quantized to 7 bits, with 3 of those bits for the fractional part. The codes examined include hypergraph product codes and lifted product QLDPC codes, as studied in [3] and [19], respectively. Specifically, the hypergraph product codes with check node degree six are B1 (882, 24, $\leq$ 24) and C2 (1922, 50, 16). The lifted product codes with check node degree eight include (442, 68, $\leq$ 10), (544, 80, $\leq$ 12), (714, 100, $\leq$ 16), and (1020, 136, $\leq$ 20). Additionally, two codes from the T family [9], T1 (126, 12, < 11) and T2 (254, 14, $<$ 17), are also analyzed.
The emulator was implemented on an AMD Virtex UltraScale+ FPGA VCU118 [23]. The evaluated architecture for BP utilized a fully parallel MS algorithm with a flooded scheduling approach from [22]. The maximum frequency for the decoders was set at 150 MHz, with two clock cycles required per iteration, equating to 13.4 ns per iteration. The number of noise generators was consistently set to $NG=40$ across all experiments, resulting in a noise generation latency ranging from 4 clock cycles for the shortest code to 26 clock cycles for the largest one. In these experiments, the thresholds of the noise generators were configured to emulate phenomenological noise.
For the shortest code, the latency of noise generation would equate to approximately 2 iterations of the MS decoder, while in the worst case, it could reach 13 iterations due to the latency introduced by the noise generators. Running all simulations at a physical error rate ranging from $10^{-3}$ to $10^{-1}$ on VCU118 took 4 hours and 48 minutes in the worst-case scenario, achieving a logical error rate of $10^{-9}$. For the (1020, 136) QLDPC code, emulating more than $10^{13}$ samples required 20 days. Faster simulations could be achieved by increasing $NG$ or using multiple boards. To replicate the same number of scenarios using the efficient software library from [24] on an Intel Core i9-14900KF [25] with 128GB of RAM, the estimated time would exceed one year. Although this result could potentially be improved by increasing parallelism or utilizing a high-performance computing (HPC) system, but even with that, the simulation would not be able to replicate the exact behavior or the hardware, which is the aspect that we will also try to exploit as we will explain in the next sections.
Beyond speed, our experiments yielded other important insights. For example, the impact of distance on the codes with degree eight (represented in purple) diminishes after a physical error rate of $10^{-5}$, where there is no aggressive error floor, but instead, a change in performance slope arises from the structural characteristics of the codes, likely constrained by the structure of the code. Additionally, we observe distinct trends between QLDPC codes with different degree distributions. The codes with degree six exhibit a significant degradation with BP, necessitating a physical error rate that is extremely low to achieve a logical error rate below $10^{-12}$, making them challenging to decode without post-processing. Conversely, the degree eight codes reach the target logical error rate before encountering a physical error rate of $10^{-4}$. It is obvious that these codes face different limitations, such as the connectivity of degree eight on the quantum side [20], but the behavior decoded looks interesting, and as we will discuss in the next section, their error correction performance can be improved just by using BP.
@@ 3 Results for BP decoders
Figure 3: Logical error rate simulations for eight QLDPC codes of degrees 6 and 8 obtained with the proposed emulator.
In addition to the previous results, we analyzed the effect of quantization noise not only to understand its impact on logical error rate performance and hardware savings in terms of area, power, and time but also because some studies have shown that a certain level of noise can enhance the convergence of BP decoders [14], [13], [10]. For this reason, we tested different quantization schemes with the previous codes to determine if we could benefit from the quantization noise.
Some of the quantization schemes we tested include 8 bits, with 4 of them fractional (q[8,4]); 7 bits, with 3 of them fractional (q[7,3]); and an extreme case of 4 bits, with 2 fractional (q[4,2]). Fig. 4 presents a comparison of the (1020,136) QLDPC code. For the sake of simplicity, we omitted the results for the other codes and quantization schemes, but similar conclusions were drawn.
As observed, the most trivial outcome is that using only 4 bits significantly degrades performance compared to using 7 or 8 bits. Additionally, the performance difference until a logical error rate of $10^{-4}$ seems reasonable compared to classical LDPC decoders; thus, 7 or 8 bits appear sufficient for message exchange [26]. However, a surprising finding is that below a logical error rate of $10^{-4}$, the scheme with fewer bits outperforms those with more bits, indicating that quantization noise may assist in the convergence of the algorithm.
Exploiting this noise that comes out naturally in hardware implementations could be beneficial and worth exploring, as opposed to other proposals that artificially introduce noise [13], [14], and [18]. This raises a critical question: Does the noise generated by different quantization schemes affect various error patterns? In other words, are we correcting the same set of errors with a larger number of bits as we do with fewer bits? This is the starting point for our diversity method proposal, described in the next section.
@@
Figure 4: Effect of various quantization schemes on the logical error rate for a QLDPC code with a check node degree of 8 obtained with the proposed emulator.
@@ 4 Diversity based on quantization noise
One of the main advantages of using the proposed emulator is its ability to store the error patterns that lead to logical errors after decoding with real hardware, especially for low logical error rates. This capability is significant because, as discussed in the previous section, having more bits does not necessarily mean there will be fewer errors. In other words, certain error patterns that a floating-point decoder cannot correct may be successfully addressed by a finite precision decoder that uses fewer bits. Therefore, capturing the patterns that cause decoding failures in real hardware is essential for effective analysis.
The first step was to fix a specific seed in the random noise generators to ensure the same scenario and perform identical simulations (with the same inputs) for the different quantization schemes. We then captured the patterns of failure and compared them offline. Our initial conclusion from this experiment was that the sets of failure patterns were different by a significant percentage, confirming our hypothesis that different quantization schemes introduce varying levels of noise that affect decoding in distinct ways. Additionally, we observed that quantization schemes using fewer bits often forced convergence, frequently resulting in incorrect codewords. This observation was crucial, as it is preferable to experience a larger number of decoding failures than to converge on incorrect codewords because otherwise, errors may go undetected.
@@
Additionally, it is important to note that the total number of bits used is less than 32 (the typical width of floating point operations). This limitation not only enhances the speed of implementation compared to software but also improves efficiency in terms of power consumption.
As an example, with the code (1020,136) and a physical error rate close to $10^{-3}$, decoders that function as post-processors are activated only $10^{-7}\%$ of the time. Among these activations, the decoder with the highest accuracy succeeds in $95\%$ of the calls, while the one with the lowest accuracy only succeeds $3.5\%$ of the time. When the physical error rate is closer to $10^{-2}$, the most accurate decoder succeeds in just $78\%$ of the cases, and the second decoder succeeds $16\%$ of the time. In this scenario, the post-processor is called upon $10^{-4}\%$ of the time.
This approach was applied to all the lifted product codes of degree eight discussed in the previous section. In Fig.5, we summarize the results, which show improvements for all codes. Additionally, some gain derived from the code distance is partially recovered in the logical error rate region below $10^{-10}$. In some instances, the gain in logical error rate exceeds one order of magnitude when the physical error rate improves.
@@
The results presented here are just a preliminary approximation; further analysis is needed to optimize these quantization schemes, which will play a crucial role when we transition to FPGA or ASIC implementations as quantum systems scale up and more efficient solutions are sought.
Figure 5: Logical error rate simulations for four QLDPC codes of degree 8 obtained with the proposed emulator and the diversity approach based on quantization noise.
@@ 5 Diversity based on BP implementations
The decoder described in the previous section operates under phenomenological noise and is associated with the hardware implementation of the algorithm. In this section, we aim to extend the idea of the diversity decoder to circuit-level noise, verifying it through a software implementation based on Stim [6] and the software library from [24]. Instead of introducing diversity through quantization noise, we introduce it via the algorithm used to implement BP, variations on the scaling factor, and modifications on the a priori information. Our approach includes providing feedback to subsequent decoders if the previous ones do not achieve convergence. Unlike other methods, such as SI or CA proposed in [9] and [10], we only modify the a priori information based on hard decisions, intentionally avoiding sorting steps or graph analysis, which are computationally expensive when the detector error model is applied. When diverse BP has a decoding failure, we apply a post-processor like LSD or OSD, but we significantly reduce the number of calls to the last. Furthermore, our method increases the degree of parallelism compared to other solutions while maintaining a similar area as a single BP decoder and minimizing latency.
The approach begins with a BP decoder with higher accuracy. In the event of a decoding failure, we activate two additional decoders: one that utilizes parameters closer to the optimal BP, and another that introduces more diversity, following a tree structure. The goal is to tackle noisy scenarios by deliberately selecting non-optimal parameters to increase the number of decodable cases, similarly to the strategy proposed in [18] for MWPM.
Once we activate the two decoders, we take the hard decision from the decoder that encountered a convergence failure and use it to adjust the a priori information for the subsequent decoder. To further enhance diversity, we apply different modification factors, denoted as $\gamma_{i}$, where $i$ indicates the decoder number, like $\gamma_{i}\times\mathbf{e^{\prime}}+(1-\gamma_{i})\times\mathbf{y}$, where $\mathbf{e^{\prime}}$ is the hard decision estimated with the decoder from the previous stage and $\mathbf{y}$ is the a priori information from the error model.
@@
    If it diverges, we run two parallel implementations of BP: one with accurate min-sum and another that is more diverse, utilizing different $\alpha$ factors.
@@
Figure 6: Logical error rate for the bicycle bivariate code (72,12,6) under circuit level noise for the proposed diversity decoders and BP and BP+OSD.
Figure 7: Logical error rate for the bicycle bivariate code (108,8,10) under circuit level noise for the proposed diversity decoders and BP and BP+OSD.
Figure 8: Logical error rate for the bicycle bivariate code (144,12,12) under circuit level noise for the proposed diversity decoders and BP and BP+OSD.
@@
We applied our proposed decoder to bivariate bicycle codes of different lengths: (72, 12, 6), (108, 8, 10), and (144, 12, 12) [20]. The results were consistent across these simulations. For the diversity-based decoding process, we utilized a BP Sum-Product algorithm with 10 iterations as the base decoder. We included two Min-Sum decoders, each with scaling factors of $\alpha=$ 0.9 and 0.75, and $\gamma_{0}=\gamma_{1}=$ 0.75, also executing 10 iterations during the second stage. In the final stage, we employed one Min-Sum decoder with a scaling factor of $\alpha=$ 0.5, $\gamma_{2}=$ 0.5 and 10 iterations, along with a combination of Min-Sum and LSD or OSD, with 2 iterations and $\gamma_{3}=$ 0.5. This setup is summarized in Fig. 9.
@@
The worst-case latency for our approach is 22 iterations of BP plus the latency of LSD or OSD.
@@
The baseline BP is executed for 100 iterations, while the benchmark BP+OSD is carried out following these 100 iterations of BP. The experimental setup introduces circuit-level noise as described in [12].
@@
In all cases analyzed, as illustrated in Figures 6, 7, and 8, the logical error rate is at least equivalent to that of BP+OSD. However, our approach requires significantly fewer calls to OSD and fewer BP iterations, as we analyze next.
@@
The results show that applying diversity decoding without LSD or OSD already improves the logical error rate by an order of magnitude for a physical error rate of 0.001 in all cases. When the post-processing is activated, this improvement is reflected in a reduction in the number of LSD or OSD activations, as summarized in Table 1. This technique results in greater savings with larger codes and demonstrates better performance with increasing physical error rates, unlike other proposed methods.
@@
Maintaining the post-processor after the diversity warrantees having at least the same performance. The reduced impact on average speedup in relation to decreasing physical error rates is attributed to the numerous cases in which BP converges within a small number of iterations. Contrarily, in higher-noise environments, the average latency is primarily influenced by LSD or OSD time.
@@
When analyzing the worst-case scenario, the latency decreases in line with the reduction in the number of OSD calls. It is well established that, in worst-case situations, the post-processing time predominates rather than the BP time. However, it is crucial to highlight that our diversity proposal results in running 70 fewer iterations than BP+OSD, leading to savings in both time and power consumption without losing accuracy.
@@
When compared to the BP+BP+OSD method presented in [12], our diversity decoder combined with OSD saves 100 iterations over the sparsified detector error model. The diversity proposal also achieves a larger performance gain compared to BP+BP without the need for two separate graphs, which is crucial from a hardware standpoint.
@@
Although this diversity-based approach needs further exploration across more code families and various configuration parameters, it already demonstrates promising gains compared to other methods.

### v2

@@ 2.5 Control layer
This communication is particularly crucial when the number of noise generators $NG$ is less than the number of required noise samples $n$. In such cases, $\left\lceil\frac{n}{NG}\right\rceil$ cycles are necessary, and the decoder may finish before the next round of samples is generated due to the early stopping criterion (see Scenario B in Fig. 2).
@@
Counterintuitively, in regions where the physical error rate is reduced, the time bottleneck can occur during noise generation. This is because the decoders require a large amount of area resources and leave insufficient space on the FPGA, even for the largest existing devices, to introduce enough noise generators to perform all the computations in parallel.
@@
We can generally categorize two scenarios: one in which the decoder consumes more clock cycles than the noise generation (Scenario A in Fig. 2), and another where the decoder is faster (Scenario B). For instance, in a fully parallel BP decoder, the dividing line between both scenarios occurs when the number of iterations required for convergence is fewer than $\left\lceil\frac{2n}{NG}\right\rceil$, assuming two clock cycles per iteration, one for computing the check nodes and another for computing the variable nodes.
@@
The signals exchanged between both finite state machines include “noise ready”, “generate noise”, and “end of decoding”. For simplicity, other control signals, such as those used to store error patterns in RAM when a logical failure occurs (for subsequent a posteriori analysis), have been omitted.
@@
Figure 3: LER simulations for eight QLDPC codes of degrees 6 and 8 obtained with the implemented emulator.
2.6 Results for BP decoders
The emulator discussed in the previous section was tested with various QLDPC codes, which are defined as ($n$, $k$, $d$), where $n$ is the number of physical qubits, $k$ is the number of logical qubits, and $d$ is the minimum distance of the code. Fig. 3 presents a small sample of eight codes with check node degrees [^3] The check node degree is the number of messages computed by a check node in a QLDPC code. of six and eight, decoded using the scaled min-sum (MS) algorithm quantized to 7 bits, with 3 of those bits for the fractional part. The codes examined include hypergraph product codes and lifted product QLDPC codes, as studied in [3] and [23], respectively  [^4] The code names B1, C2, T1, and T2 are maintained from references [3, 23] to help readers identify the corresponding construction methods in the original works.. Specifically, the hypergraph product codes with check node degree six are B1 (882, 24, $\leq$ 24) and C2 (1922, 50, 16). The lifted product codes with check node degree eight include (442, 68, $\leq$ 10), (544, 80, $\leq$ 12), (714, 100, $\leq$ 16), and (1020, 136, $\leq$ 20). Additionally, two codes from [12], T1 (126, 12, < 11) and T2 (254, 14, << 17), are also analyzed.
The emulator was implemented on an AMD Virtex UltraScale+ FPGA VCU118 [26], but it is worth noting that it can be ported to any FPGA with sufficient logic resources. The evaluated architecture for BP utilized a fully parallel MS algorithm with a flooded scheduling approach from [4]. The maximum frequency for the decoders was set at 150 MHz, with two clock cycles required per iteration, equating to 13.3 ns per iteration. The number of noise generators was consistently set to $NG=40$ across all experiments, resulting in a noise generation latency ranging from 4 clock cycles for the shortest code (T1, $\left\lceil\frac{126}{40}\right\rceil=4$) to 49 clock cycles for the largest one (C2, $\left\lceil\frac{1922}{40}\right\rceil=49$). In these experiments, the thresholds of the noise generators were configured to emulate a code-capacity model.
@@ 2.6 Results for BP decoders
Figure 4: Effect of various quantization schemes on the LER for a QLDPC code with a check node degree of 8 obtained with the proposed emulator.
For the shortest code, the latency of noise generation would equate to approximately 2 iterations of the MS decoder, while in the worst case, it could reach 13 iterations due to the latency introduced by the noise generators. Running all simulations at a physical error rate ranging from $10^{-3}$ to $10^{-1}$ on VCU118 took 4 hours and 48 minutes in the worst-case scenario, achieving a LER of $10^{-9}$. For the (1020, 136) QLDPC code, emulating more than $10^{13}$ samples required 20 days. Faster simulations could be achieved by increasing $NG$ or using multiple boards. To replicate the same number of scenarios using the efficient software library from [27] on an Intel Core i9-14900KF [28] with 128GB of RAM, the estimated time would exceed one year. Although this result could potentially be improved by increasing parallelism or utilizing a high-performance computing (HPC) system, even with that, the simulation would not be able to replicate the exact behavior of the hardware, which is the aspect that we will also try to exploit, as we will explain in the next sections.
Beyond speed, our experiments yielded other important insights. For example, the impact of distance on the codes with degree eight (represented in purple) diminishes after a LER of $10^{-5}$, where there is no aggressive error floor. Additionally, we observe distinct trends between QLDPC codes with different degree distributions. The codes with degree six exhibit a significant degradation with BP, necessitating a physical error rate that is extremely low to achieve a LER below $10^{-12}$, making them challenging to decode without post-processing. Conversely, the degree eight codes reach the target LER before encountering a physical error rate of $10^{-4}$. It is obvious that these codes face different limitations, such as the connectivity of degree eight on the quantum side [24], but the behavior decoded looks interesting, and as we will discuss in the next section, their error correction performance can be improved just by using BP.
In addition to the previous results, we analyzed the effect of quantization noise not only to understand its impact on LER performance and hardware savings in terms of area, power, and time but also because some studies have shown that a certain level of noise can enhance the convergence of BP decoders [18], [17], [13]. For this reason, we tested different quantization schemes for the messages computed in BP to determine if we could benefit from the quantization noise.
Some of the quantization schemes we tested include 8 bits, with 4 of them to represent the fractional part of the messages (q[8,4]); 7 bits, with 3 of them fractional (q[7,3]); and an extreme case of 4 bits, with 2 fractional (q[4,2]). Fig. 4 presents a comparison of the LER obtained with a BP decoder for the (1020,136) QLDPC code applying different quantization schemes. For the sake of simplicity, we omitted the results for the other codes and quantization schemes, but similar conclusions were drawn.
As observed, the most trivial outcome is that using only 4 bits significantly degrades performance compared to using 7 or 8 bits. Additionally, the performance difference until a LER of $10^{-4}$ seems reasonable; thus, 7 or 8 bits appear sufficient for message exchange between check nodes and variable nodes [29]. However, a surprising finding is that below a LER of $10^{-4}$, the scheme with fewer bits outperforms those with more bits, indicating that quantization noise may assist in the convergence of the algorithm. The observation that lower bit-width quantization can outperform higher precision at low logical error rates (below $10^{-4}$) is not new to the decoding literature. In classical LDPC decoding [30], [31], [32], [33], prior works have demonstrated that coarse quantization limits message magnitudes, suppresses harmful feedback loops, and mitigates trapping-set effects, especially in the error-floor regime. However, quantum LDPC codes differ from their classical counterparts due to degeneracy and strong correlations in their graph structure [3], which invalidate traditional analytic methods such as density evolution. As a result, hardware-level emulation becomes essential for accurately assessing decoding performance under finite-precision constraints. In this context, quantization introduces structured noise that disrupts feedback loops and prevents cycles from reinforcing erroneous messages, which can improve decoding accuracy. While further theoretical development is needed to predict the most effective quantization schemes without exhaustive emulation, this phenomenon highlights the need for innovative approaches to analyze quantum LDPC decoders, where classical methods fall short. In the meantime, emulation remains the only reliable way to verify the behavior of quantum LDPC decoders under finite-precision constraints.
Exploiting this noise that comes out naturally in hardware implementations could be beneficial and worth exploring, as opposed to other proposals that artificially introduce noise [17], [18], and [22]. This raises a critical question: Do different quantization schemes in the decoder lead to the correction of a different set of error patterns due to biases introduced by quantization noise? In other words, are we correcting the same set of error patterns with a larger number of bits as we do with fewer bits, or are we just correcting more errors but of a different set? This is the starting point for our diversity method proposal, described in the next section.
@@ 3 Diversity based on quantization noise
One of the main advantages of using the implemented emulator is its ability to store the error patterns that lead to logical errors after decoding with real hardware, especially for low LERs. This capability is significant because, as discussed in the previous section, having more bits does not necessarily mean there will be fewer errors. In other words, certain error patterns that a floating-point decoder cannot correct may be successfully addressed by a finite precision decoder that uses fewer bits. Therefore, capturing the patterns that cause decoding failures in real hardware is essential for effective analysis.
The first step was to fix a specific seed in the random noise generators to ensure the same scenario and perform identical simulations (with the same inputs) for the different quantization schemes. We then captured the patterns of failure and analyzed their differences when decoders with different quantization schemes were applied offline. Our initial conclusion from this experiment was that the sets of failure patterns were different by a significant percentage, confirming our hypothesis that different quantization schemes introduce varying levels of noise that affect decoding in distinct ways. Additionally, we observed that quantization schemes using fewer bits often forced convergence, frequently resulting in incorrect codewords. This observation was fundamental, as it is preferable to experience a larger number of decoding failures than to converge on incorrect codewords because otherwise, errors may go undetected.
@@
Additionally, it is important to note that the total number of bits used is less than 32 (the typical width of single-precision floating-point operations on a general-purpose CPU). This reduction not only enhances the speed of implementation compared to software but also improves efficiency in terms of power consumption.
As an example, with the code (1020,136) and a physical error rate close to $10^{-3}$, decoders that function as post-processors (MS with q[8,4], q[4,2], and q[3,1]) are activated only $10^{-7}\%$ of the time. Among these activations, the decoder with the highest accuracy succeeds in $95\%$ of the calls, while the one with the lowest accuracy only succeeds $3.5\%$ of the time. When the physical error rate is closer to $10^{-2}$, the most accurate decoder succeeds in just $78\%$ of the cases, and the second decoder succeeds $16\%$ of the time. In this scenario, the post-processor is called upon $10^{-4}\%$ of the time.
This approach was applied to all the lifted product codes of degree eight discussed in the previous section. In Fig.5, we summarize the results, which show improvements for all codes. Additionally, some gain derived from the code distance is partially recovered in the LER region below $10^{-10}$. In some instances, the gain in LER exceeds one order of magnitude when the physical error rate improves.
Figure 5: LER simulations for four QLDPC codes of degree 8 obtained with the proposed emulator and the diversity approach based on quantization noise.
@@
The results presented here are just a preliminary approximation; further analysis is needed to optimize these quantization schemes, which will play an essential role when we transition to FPGA or ASIC implementations as quantum systems scale up and more efficient solutions are sought.
@@ 4 Diversity based on BP implementations
The decoder described in the previous section operates under code-capacity model. In this section, we aim to extend the idea of the diversity decoder to circuit-level noise, verifying it through a software implementation based on Stim [9] and the software library from [27]. Instead of introducing diversity through quantization noise, we achieve it by varying the scaling factor, modifying the a priori information, and combining BP decoders with different update rules. Our approach includes providing feedback to subsequent decoders if the previous ones do not achieve convergence. Unlike other methods, such as stabilizer inactivation or check-agnosia proposed in [12] and [13], we only modify the a priori information based on the binary outputs of the previous decoder (the hard decisions of the estimated error vector), intentionally avoiding sorting steps or graph analysis, which are computationally expensive when the detector error model is applied. When diverse BP has a decoding failure, we apply a post-processor like LSD or OSD, but we significantly reduce the number of calls to the latter. Furthermore, our method increases the degree of parallelism compared to other solutions that are based on a concatenation of sequential decoders, while maintaining similar logical resources (FPGA area) as a single BP decoder and minimizing latency.
The approach begins with a BP decoder with higher accuracy. In the event of a decoding failure, we activate two additional decoders: one that utilizes a scaling factor that provides a LER closer to the optimal BP, and another that introduces more diversity, following a tree structure. The goal is to tackle noisy scenarios by deliberately selecting non-optimal parameters to increase the number of decodable cases, similarly to the strategy proposed in [22] for MWPM.
Once we activate the two decoders, we take the hard decision from the decoder that encountered a convergence failure and use it to adjust the a priori information for the subsequent decoder.
We define the a priori information vector $\mathbf{y}$ as the log-likelihood ratio, $y_{j}=\log\left(\frac{\mathbb{P}(e_{j}=0)}{\mathbb{P}(e_{j}=1)}\right)$ with $j\in\{1,\dots,n\}$ and $\mathbf{e}\in\mathbb{F}_{2}^{n}$ denote the binary error vector. The error vector estimated by the decoder is defined as $\mathbf{e^{\prime}}\in\mathbb{F}_{2}^{n}$, which is the binary vector based on the hard decision of the computed messages.
To further enhance diversity, we apply different modification factors, denoted as $\gamma_{i}$, where $i$ indicates the decoder number, like $\gamma_{i}\times\mathbf{e^{\prime}}+(1-\gamma_{i})\times\mathbf{y}$.
@@
    If it diverges, we run two parallel implementations of BP: one with accurate min-sum and another that is more diverse, utilizing different scaling factors for the exchanged messages, $\alpha$.
@@
Figure 7: LER for the bicycle bivariate code (72,12,6) under circuit level noise for the proposed diversity decoders and BP and BP+OSD.
Figure 8: LER for the bicycle bivariate code (108,8,10) under circuit level noise for the proposed diversity decoders and BP and BP+OSD.
Figure 9: LER for the bicycle bivariate code (144,12,12) under circuit level noise for the proposed diversity decoders and BP and BP+OSD.
@@
We applied our proposed decoder to bivariate bicycle codes of different lengths: (72, 12, 6), (108, 8, 10), and (144, 12, 12) [24]. The results were consistent across these simulations. For the diversity-based decoding process, we utilized a BP Sum-Product algorithm with 10 iterations as the base decoder. We included two Min-Sum decoders, each with scaling factors of $\alpha=$ 0.9 and 0.75, and $\gamma_{0}=\gamma_{1}=$ 0.75, also executing 10 iterations during the second stage. In the final stage, we employed one Min-Sum decoder with a scaling factor of $\alpha=$ 0.5, $\gamma_{2}=$ 0.5 and 10 iterations, along with a combination of Min-Sum and LSD or OSD, with 2 iterations and $\gamma_{3}=$ 0.5. The values of the scaling factors and coefficients used in the diversity decoder are chosen to enable a hardware-friendly implementation rather than arbitrarily. These values allow the decoder to avoid hardware multipliers by implementing the products using simple bit shifts and additions. This design minimizes hardware resource usage and the latency while maintaining effective scaling. This setup is summarized in Fig. 6.
@@
The worst-case latency for our approach is 22 iterations of BP plus the latency of LSD or OSD, as 30 BP iterations of the parallel schedule is less restrictive.
@@
The baseline BP is executed for 100 iterations, while the benchmark BP+OSD is carried out following these 100 iterations of BP. All the simulations are performed under circuit-level noise as described in [15] and for the LSD implementation, we refer to the authors’ library [27].
@@
In all cases analyzed, as illustrated in Figures 7, 8, and 9, the LER of the proposed decoder is at least equivalent to that of BP+OSD. However, our approach requires significantly fewer calls to OSD and fewer BP iterations, as we analyze next.
@@
The results show that applying diversity decoding without LSD or OSD already improves the LER by an order of magnitude compared to BP for a physical error rate of 0.001 in all cases. When the post-processing is activated, this improvement is reflected in a reduction in the number of LSD or OSD activations, as summarized in Table 1. This technique results in greater savings with larger codes and demonstrates better performance with increasing physical error rates.
@@
Maintaining the post-processor after the diversity warrants having at least the same performance. The reduced impact on average speedup in relation to decreasing physical error rates is attributed to the numerous cases in which BP converges within a small number of iterations. Contrarily, in higher-noise environments, the average latency is primarily influenced by LSD or OSD time.
@@
When analyzing the worst-case scenario, the latency decreases in line with the reduction in the number of OSD calls. It is well established that, in worst-case situations, the post-processing time predominates rather than the BP time. However, it is important to highlight that our diversity proposal results in running 70 fewer iterations than BP+OSD, leading to savings in both time and power consumption without losing accuracy.
@@
When compared to the BP+BP+OSD method presented in [15], our diversity decoder combined with OSD saves 100 iterations over the sparsified detector error model. The diversity proposal also achieves a larger performance gain compared to BP+BP without the need for two separate graphs, which can be important from a hardware standpoint.
@@
In comparison with recent competitive decoders, such as Relay-BP [5], our proposed diversity method offers a different trade-off. For the [[144,12,12]] QLDPC code, Relay-BP, which uses 301 relay legs, achieves a significant improvement in decoding performance, reducing the logical error rate by nearly an order of magnitude compared to BP+OSD. In contrast, our diversity approach, using only four BP decoders and a post-processor, results in a more modest improvement over BP+OSD, but it requires fewer decoders and still provides a 2.2x speedup in the worst case. While our solution does not improve BP+OSD as significantly as Relay-BP, it offers a more compact approach that reduces the number of decoding instances needed, making it a promising option for scenarios where minimizing computational resources is critical. Furthermore, combining our diversity approach with other methods such as Relay-BP, or SI could lead to a “sweet spot” that balances decoding performance and computational efficiency, providing a more flexible solution for different application needs and constraints.
Although this diversity-based approach requires further exploration across more code families and configuration parameters, the current results suggest promising gains, either standalone or in combination with other methods. A fully normalized comparison in terms of hardware cost and resource usage is left for future work.